A car with mass 1800 kg traveling south at a speed of 35 m/s collides at an intersection with a 2500 kg van traveling west at a speed of 40m/s. Find the magnitude and direction of the velocity of the wreckage after the collision, assuming that the vehicles undergo a perfectly inelastic collision and assuming that friction between the vehicles and the road can be neglected.



1
Expert's answer
2020-11-17T07:08:57-0500


Let's write down the momentum conservation law:


"m_1\mathbf{v}_1 + m_2\mathbf{v}_2 = (m_1+m_2)\mathbf{v}'"

where "m_1 = 1800kg", "v_1 = 35m\/s", "m_2 = 2500kg", and "v_2 = 40m\/s". "\mathbf{v}'" is their velocity after the collision. Its magnitude is:


"v' = \dfrac{|m_1\mathbf{v}_1 + m_2\mathbf{v}_2|}{m_1 + m_2}"

Since "m_1\mathbf{v}_1" and "m_2\mathbf{v}_2" form right triangle, the magnitude of their sum will be:


"|m_1\mathbf{v}_1 + m_2\mathbf{v}_2| = \sqrt{(m_1v_1)^2 + (m_2v_2)^2}"

Thus, obtain:


"v' = \dfrac{\sqrt{(m_1v_1)^2 + (m_2v_2)^2}}{m_1 + m_2}\\
v' = \dfrac{\sqrt{(1800\cdot 35)^2 + (2500\cdot 40)^2}}{1800 + 2500} \approx 27.49\space m/s"

The direction is shown in the figure.